# Describe-and-Dissect: Interpreting Neurons in Vision Networks with Language Models

### v1

@@ 4.1 Qualitative evaluation
We qualitatively analyze results of randomly selected neurons from various layers of ResNet-50, ResNet-18, and ViT-B-16, displayed in Figure 1 in the main text, and Figures 7, 8, 9, 10, and 11 in Appendix. We use the union of the ImageNet validation dataset and Broden as $\mathcal{D}_{probe}$ and compare to Network Dissection (Bau et al., 2017), MILAN (Hernandez et al., 2022), and CLIP-dissect (Oikarinen & Weng, 2023) as baselines. Labels for each method are color coded by whether we believe they are accurate, somewhat correct, or vague/imprecise. Compared to baseline models, we observe that DnD captures higher level concepts in a more semantically coherent manner. Specifically, methods such as CLIP-dissect and Network Dissection have limited expressability due to the use of restricted concept sets while MILAN produces labels confined to lower level concepts.
@@ 4.2 Quantitative evaluation
4.2.1 Final layer evaluation
Here we follow (Oikarinen & Weng, 2023) to quantitatively analyze description quality on the last layer neurons, which have known ground truth labels (i.e. class name) to allow us to evaluate the quality of neuron descriptions automatically. In this evaluation, we focus on comparison with MILAN (Hernandez et al., 2022), as it is the other generative contemporary work in the baselines. Network Dissection (Bau et al., 2017) and CLIP-Dissect (Oikarinen & Weng, 2023) are not included in this comparison because these methods have concept sets where the ”ground truth” class or other similar concepts can be included, giving them an unfair advantage to the methods without concept sets like MILAN and DnD. We reported the results for all of the neurons of ResNet-50’s final fully-connected layer in Table 2. Our results show that DnD outperforms MILAN, having a greater average CLIP cosine similarity by 0.0518, a greater average mpnet cosine similarity by 0.18, and a greater average BERTScore by 0.008. On average DnD’s labels are significantly closer to the ground truths than MILAN’s.
@@ 4.2.1 Final layer evaluation
4.2.2 MILANNOTATIONS evaluation
Though evaluation on hidden layers of deep vision networks can prove quite challenging as they lack ”ground truth” labels, one resource to perform such a task is the MILANNOTATIONS dataset (Hernandez et al., 2022), which collected annotated labels to serve as ground truth neuron explanations for a network. We perform quantitative evaluation by calculating the textual similarity between a method’s label and the corresponding MILANNOTATIONS. However, we found that this dataset is too noisy and not suitable for a reliable evaluation. Our analysis in section A.3 found that if every neuron is described with the same constant concept: ‘depictions‘, it will achieve better results than any explanation method on this dataset, but this is not a useful nor meaningful description. We hypothesize this is due to high levels on noise and interannotator disagreement on the dataset, leading to very generic descriptions scoring highly. We found that the average textual similarity between each neuron’s three MILANNOTATIONS is relatively low for annotations that are supposed to be describing the same concept. Thus, this dataset is unreliable to serve as ground truths and can’t be relied on for comparing different methods.
@@ 4.3 Crowdsourced experiment
Table 3: Results for individual layers of ResNet-50. We observe that DnD is the best method across all layers in ResNet-50. Our concepts achieved a mean rating of over 4.0 for every layer and were 2$\times$ as likely to be selected as the best concept when compared to baseline methods.
Method
Metric Layer Network Dissection MILAN CLIP-Dissect DnD (Ours)
Mean Rating Layer 1 $3.41\pm 0.058$ $3.41\pm 0.060$ $3.63\pm 0.057$ 4.16 $\pm$ 0.041
Layer 2 $3.14\pm 0.067$ $3.12\pm 0.064$ $3.55\pm 0.057$ 4.07 $\pm$ 0.048
Layer 3 $3.04\pm 0.066$ $2.96\pm 0.066$ $3.66\pm 0.055$ 4.14 $\pm$ 0.042
Layer 4 $2.97\pm 0.066$ $3.34\pm 0.061$ $3.82\pm 0.054$ 4.21 $\pm$ 0.044
% time Layer 1 13.18% 14.32% 22.50% 50.00%
selected Layer 2 15.27% 12.41% 19.33% 52.98%
as best Layer 3 11.82% 12.73% 25.00% 50.45%
answer Layer 4 10.56% 13.71% 25.62% 50.11%
Table 4: Results for individual layers of ResNet-18. DnD performs significantly better across every layer in ResNet-18 when compared to the baseline methods. Additionally, our concepts are selected over more than twice as often than any other method.
Method
Metric Layer Network Dissection MILAN CLIP-Dissect DnD (Ours)
Mean Rating Layer 1 $3.56\pm 0.102$ $3.08\pm 0.115$ $3.39\pm 0.122$ 4.18 $\pm$ 0.081
Layer 2 $3.03\pm 0.107$ $3.00\pm 0.117$ $3.45\pm 0.106$ 4.22 $\pm$ 0.082
Layer 3 $3.14\pm 0.120$ $2.98\pm 0.114$ $3.46\pm 0.108$ 4.08 $\pm$ 0.079
Layer 4 $3.60\pm 0.110$ $3.49\pm 0.116$ $3.77\pm 0.100$ 4.10 $\pm$ 0.081
% time Layer 1 15.09% 11.32% 19.81% 53.77%
selected Layer 2 6.31% 11.71% 19.82% 62.16%
as best Layer 3 14.29% 12.38% 15.24% 58.10%
answer Layer 4 15.09% 17.92% 22.64% 44.34%
@@
We dissected both a ResNet-50 network pretrained on Imagenet-1K and ResNet-18 trained on Places365, using the union of ImageNet validation dataset and Broden (Bau et al., 2017) as our probing dataset. For both models we evaluated 4 of the intermediate layers (end of each residual block), with 200 randomly chosen neurons per layer for ResNet50 and 50 per layer for ResNet-18. Each neurons description was evaluated by 3 different workers.
@@
The full task interface and additional experiment details are available in Appendix A.1.5. Workers were presented with the top 10 highest activating images of a neuron followed by four separate descriptions; each description corresponds to a label produced by one of the four methods compared. The descriptions are rated on a 1-5 scale, where a rating of 1 represents that the user ”strongly disagrees” with the given description, and a rating of 5 represents that the user ”strongly agrees” with the given description. Additionally, we ask workers to select the description that best represents the 10 highly activating images presented.
@@
Results. Table 3 and Table 4 shows the results of a large scale human evaluation study conducted on Amazon Mechanical Turk (AMT). Looking at ”time selected as best” as the comparison metric, our results show that DnD performs over 2$\times$ better than all baseline methods when dissecting ResNet-50 or ResNet-18, being selected the best of the four up to 62.16% of the time. In terms of mean rating, our method achieves an average label rating over 4.1 for both dissected models, whereas the average rating for the second best method CLIP-Dissect is only 3.67 on ResNet-50 and 3.52 on ResNet-18. Our method also significantly outperforms MILAN’s generative labels, which averaged below 3.3 for both target models. In conclusion we have shown that our method significantly outperforms existing methods in crowdsourced evaluation, and does this consistently across different models and layers.
@@
4.4 Use case
To showcase a potential use case for neuron descriptions (and provide another way to quantitatively compare explanation methods), we experimented with using neuron descriptions to find a good classifier for a class missing from the training set. Our setup was as follows: we explained all neurons in layer4 of ResNet-50 (ImageNet) using different methods. We then wanted to find neurons in this layer that could serve as the best classifiers for an unseen class, specifically the classes in CIFAR-10 and CIFAR-100 datasets. Note there is some overlap between these and ImageNet classes, but CIFAR classes are typically much more broad. To find a neuron to serve as a classifier, we found the neuron whose description was closest to the CIFAR class name in a text embedding space (ensemble of the CLIP ViT-B/16 and mpnet text encoders used in this paper). We then measured how well that neuron (its average activation) performs as a single class classifier on the CIFAR validation dataset, measured by area under ROC curve. For cases where multiple neurons shared the closest description, we averaged the performance of all neurons with that description.
Results are shown in Table 5. We can see DnD performs quite well, reaching AUROC values around 0.74, while MILAN performs much worse. We believe this may be because MILAN descriptions are very generic (likely caused by noisy dataset, see Appendix A.3), which makes it hard to find a classifier for a specific class. We believe this is a good measure of explanation quality, as different methods are dissecting the same network, and even if no neurons exist that can directly detect a class, a better method should find a closer approximation.
MILAN DnD (Ours)
CIFAR10 0.5906 0.7036
CIFAR100 0.6514 0.7396
4.5 Ablation Studies for each step in DnD pipeline
For further insight, we conduct comprehensive ablation studies found in Section A.4 analyzing the importance of each step in the DnD pipeline. In Appendix A.4.1, we perform an attention cropping ablation study on DnD Step 1 to qualitatively show that attention cropping effectively eliminates spurious concepts shown in Figure 15. We also perform an image captioning ablation study on DnD Step 2A, matching labels from a fixed concept set to highly activating images rather than generating image captions, and showcase in Appendix A.4.2, Table 8, and Figures 16 and 17 that generative image captions outperform a fixed concept set. We further perform an additional ablation study on DnD Step 2A for our image-to-text model, BLIP (Li et al., 2022), and both qualitatively and quantitatively compare it with BLIP2 (Li et al., 2023) in Appendix A.4.3, Table 9, and Figures 18 and 19 to show that they perform quite similarly and in certain important scenarios, BLIP outperforms BLIP2 for our purposes. In Appendix A.4.4 and Figure 20 we study the effects of DnD Step 2B (GPT concept summarization) and qualitatively show how it is necessary to encapsulate the broad concept of a neuron. Finally, we conduct an ablation study on DnD Step 3 (Best Concept Selection) in Appendix A.4.5 and show through Table 10 and Figure 3 that this step consistently improves label quality by eliminating vague/imprecise labels.
@@ 5 Conclusions
In this paper, we presented Describe-and-Dissect (DnD), a novel method for automatically labeling the functionality of deep vision neurons without the need for labeled training data or a provided concept set. We accomplish this through three important steps including probing set augmentation, candidate concept generation through off-the-shelf general purpose models, and best concept selection with carefully designed scoring functions. Through extensive qualitative and quantitative analysis, we show that DnD outperforms prior work by providing significantly higher quality neuron descriptions, while being more general and flexible than existing methods.
Impact Statement
Our work aims to improve our understanding of machine learning models, which we expect to have positive societal impact as more understanding will decrease the chance of harm from deploying unsafe or unreliable models. However, one potential downside may be that if explanations look convincing but are not reliable enough, they may cause a user to place unwarranted trust on an unreliable model.

### v2

@@ 4.1 Qualitative evaluation
We qualitatively analyze results of randomly selected neurons from various layers of ResNet-50, ResNet-18, and ViT-B-16. Sample results are displayed in Figure 1 and Figures 9, 10, 11, 12, 13, 15, and 16 in the Appendix. We use the union of the ImageNet validation dataset and Broden as $\mathcal{D}_{probe}$ and compare to Network Dissection (Bau et al., 2017), MILAN (Hernandez et al., 2022), and CLIP-dissect (Oikarinen & Weng, 2023) as baselines. Labels for each method are color coded by whether we believe they are accurate, somewhat correct, or vague/imprecise. Compared to baseline models, we observe that DnD captures higher level concepts in a more semantically coherent manner. Specifically, methods such as CLIP-dissect and Network Dissection have limited expressability due to the use of restricted concept sets while MILAN produces labels confined to lower level concepts. Additionally, we find that DnD can express multiple concepts within a single label owing to its generative nature.
@@ 4.2 Quantitative Evaluation
Final layer evaluation.
Here we follow Oikarinen & Weng (2023) to quantitatively analyze description quality on the last layer neurons, which have known ground truth labels (i.e. class name) to allow us to evaluate the quality of neuron descriptions automatically. To evaluate on accuracy, we measure the average CLIP and mpnet cosine similarity between ground-truth labels and generated descriptions on all output neurons of ResNet-50’s final fully-connected layer. BERTScore metric measures the semantic interpretability of DnD descriptions. We focus this study on comparison with MILAN (Hernandez et al., 2022) as it is the other generative contemporary work in the baselines. Network Dissection (Bau et al., 2017) and CLIP-Dissect (Oikarinen & Weng, 2023) are not included in this comparison because these methods have concept sets where the "ground truth" class or other similar concepts can be included, giving them an unfair advantage to the methods without concept sets like MILAN and DnD. We reported the results for all of the neurons of ResNet-50’s final fully-connected layer in Table 2. Our results show that DnD outperforms MILAN, producing labels that are significantly closer to the ground truths than MILAN’s.
@@ 4.3 Crowdsourced experiment
Table 3: Averaged AMT results across four layers in ResNet-50 and ResNet-18. We can see DnD outperforms existing methods on both ResNet-50 (ImageNet) and ResNet-18 (Places365). Our descriptions are consistently rated the highest and selected as the best more than twice as often as the best baseline.
Metric / Method ResNet-50 ResNet-18
NetDissect MILAN CLIP-Dissect DnD (Ours) NetDissect MILAN CLIP-Dissect DnD (Ours)
Mean Rating 3.14±$\pm{0.032}$ 3.21±$\pm{0.032}$ 3.67±$\pm{0.028}$ 4.15±$\pm{0.022}$ 3.33±$\pm{0.056}$ 3.14±$\pm{0.059}$ 3.52±$\pm{0.055}$ 4.14±$\pm{0.041}$
Selected as Best 12.71% 13.29% 23.11% 50.89% 12.62 13.32% 19.39% 54.67%
@@
We dissected both a ResNet-50 network pretrained on Imagenet-1K and ResNet-18 trained on Places365, using the union of ImageNet validation dataset and Broden (Bau et al., 2017) as our probing dataset. For both models we evaluated 4 of the intermediate layers (end of each residual block), with 200 randomly chosen neurons per layer for ResNet50 and 50 per layer for ResNet-18. Each neurons description was evaluated by 3 different workers. In total, 3000 human ratings were conducted, 2400 evaluations on ResNet-50 and 600 evaluations on ResNet-18.
@@
The full task interface and additional experiment details are available in Appendix A.1.6. Workers were presented with the top 10 highest activating images of a neuron followed by four separate descriptions; each description corresponds to a label produced by one of the four methods compared. The descriptions are rated on a 1-5 scale, where a rating of 1 represents that the user "strongly disagrees" with the given description, and a rating of 5 represents that the user "strongly agrees" with the given description. Additionally, we ask workers to select the description that best represents the 10 highly activating images presented. For these highly activating images, we used the images calculated by our method. As our probing dataset is a superset of the image sets used by prior methods, we believe our model is the most accurate for determining images to visualize since the probing dataset encapsulates the most concepts.
@@
Results. Table 3 shows the results of a large scale human evaluation study conducted on Amazon Mechanical Turk (AMT). Looking at "time selected as best" as the comparison metric, our results show that DnD performs over 2$\times$ better than all baseline methods when dissecting ResNet-50 or ResNet-18, being selected the best of the four up to 54.67% of the time. In terms of mean rating, our method achieves an average label rating over 4.1 for both dissected models, whereas the average rating for the second best method, CLIP-Dissect, is only 3.67 on ResNet-50 and 3.52 on ResNet-18. Our method also significantly outperforms MILAN’s generative labels, which averaged below 3.3 for both target models. Our method significantly outperforms existing methods in crowdsourced evaluation, and does this consistently across different models and layers.
4.3.1 MILANNOTATIONS evaluation
Though evaluation on hidden layers of deep vision networks can prove quite challenging as they lack "ground truth" labels, one resource to perform such task is the MILANNOTATIONS dataset (Hernandez et al., 2022), which collects annotated labels to serve as ground truth neuron explanations. We perform quantitative evaluation by calculating the textual similarity between a method’s label and the corresponding MILANNOTATIONS. Our analysis in Section A.3 found that if every neuron is described with the same constant concept "depictions", it will achieve better results than any explanation on the dataset, but this is not a useful nor meaningful description. We hypothesize this is due to high levels of noise and interannotator disagreement, leading to low textual similarity between descriptions and generic descriptions scoring highly. We conclude that this dataset is unreliable to serve as ground truths for comparing different methods.
4.4 Ablation Studies
4.4.1 DnD with fixed concept set
To analyze the importance of using a generative image-to-text model, we explore instead utilizing fixed concept sets with CLIP (Radford et al., 2021) to generate descriptions for each image instead of BLIP, while the rest of the pipeline remains unchanged (i.e. using GPT to summarize etc). For the experiment, we use CLIP-ViT-B/16, where we define $L(\cdot)$ and $E(\cdot)$ as text and image encoders respectively. From the initial concept set $\mathcal{S}=\{t_{1},t_{2},...\}$, the best concept for image $I_{m}$ is defined as $t_{l}$, where $l=\text{argmax}_{i}(L(t_{i})\cdot E(I_{m})^{\top})$. Following CLIP-dissect (Oikarinen & Weng, 2023), we use $\mathcal{S}$ = 20k ($20,000$ most common English words) [^2] Source: https://github.com/first20hours/google-10000-english/blob/master/20k.txt and $\mathcal{D}_{probe}=\text{ImageNet}\cup\text{Broden}$.
To compare the performance, following Oikarinen & Weng (2023), we use our model to describe the final layer neurons of ResNet-50 (where we know their ground truth role) and compare description similarity to the class name that neuron is detecting, as discussed in Section 4.2. Results in Table 4 show that both methods perform similarly on the FC layer. In intermediate layers, we notice that single word concept captions from 20k significantly limit the expressiveness of DnD, suggesting generative image descriptions is important for our overall performance. Qualitative examples and notable failure cases of CLIP descriptions can be found under Appendix A.4.2.
Table 4: Mean FC Layer Similarity of CLIP Captioning. Utilizing a fixed concept set to caption activating images via CLIP (Radford et al., 2021), we compute the mean cosine similarity across fully connected layers of RN50. We find the performance of DnD w/ CLIP Captioning is slightly worse than BLIP generative caption.
Metric / Methods DnD (Ours) DnD w/ CLIP Captioning % Decline
CLIP cos 0.7598 0.7583 0.197%
mpnet cos 0.4588 0.4465 2.681%
BERTScore 0.8286 0.8262 0.290%
4.4.2 Effects of Concept Selection
We use 50 randomly chosen neurons from each of the 4 layers of ResNet-50 to conducted an ablation study on the impact of Best Concept Selection (Step 3) on the pipeline. Evaluations are conducted on the first candidate concept. Each neuron was evaluated twice yielding a total of 400 human ratings. Table 5 shows the effect of Best Concept Selection on the overall accuracy of DnD. We can see DnD performance is already high without Best Concept Selection, but Concept Selection further improves the quality of selected labels in Layer 2 through Layer 4, while having the same performance on Layer 1. One potential explanation is due to Layer 1 detecting more limited low level concepts–there is less variance in candidate descriptions identified in Concept Generation (Step 2), resulting in similar ratings across the set of candidate concepts $T$. We can see some individual examples of the improvement Concept Selection provides in Figure 3, with the new labels yielding more specific and accurate descriptions of the neuron. For example Layer 2 Neuron 312 becomes more specific colorful festive settings instead of generic Visual Elements.
@@ 4.4.2 Effects of Concept Selection
Method / Layer Layer 1 Layer 2 Layer 3 Layer 4 All Layers
DnD (w/o Best Concept Selection) 3.54 3.77 4.00 4.02 3.84
DnD (full pipeline) 3.54 4.00 4.24 4.13 3.97
5 Use Case: Land-Cover Prediction
One important role of interpretability tools is the ability to create real-world impact. In this section, we study applications in sustainability and climate change by applying DnD to the task of land cover prediction–a critical factor for managing water resources, conserving biodiversity, planning sustainable urban development, and mitigating climate change effects. DnD’s concept descriptions for neurons not only improve the model performance, but also identify spurious correlations, suggesting a critical role of interpretability techniques in ensuring reliability and building trust in AI systems.
Setup. We evaluate our framework on two classification models. Tile2Vec (Jean et al., 2019) utilizes a modified ResNet-18 backbone trained to minimize triplet loss between anchor, neighbor, and distant land tiles from the NAIP dataset (Claire Boryan & Craig, 2011). Following Jean et al. (2019), we train a random forest classifier on ResNet-18 embeddings to evaluate accuracy on 27 subclasses of Cropland Data Layer (CDL) labels. We also evaluate a ResNet-50 model trained on labeled EuroSAT images (Helber et al., 2019) with 10 land cover classes. We experiment on two probing datasets. NAIP is an aerial imagery dataset updated annually by the USDA. The dataset is composed of crop cover data with annotated ground truth masks from CDL labels. However, due to the fine-grained nature of the 27 subclasses, we also categorize each subclass into six broad superclasses to improve our understanding of the results: 1. Planted/Cultivated, 2. Herbaceous/Shrubland, 3. Urban/Suburban, 4. Barren, 5. Forest, 6. Water/wetlands. We use NAIP for all experiments on the Tile2Vec ResNet-18 model. On EuroSAT ResNet-50, we use the union of EuroSAT $\cup$ ImageNet $\cup$ Broden datasets as the probing dataset.
5.1 Locating Conceptual Groupings
Figure 4: Layer 2 Concept Profile. We cluster neurons with similar concepts and categorize them into 6 NAIP superclasses. Interpretable concepts have more neurons associated with them. Some superclasses do not appear due to dataset bias or intrinsic similarities between classes.
We identify neuron clusters detecting similar concepts across Layer 2 of Tile2Vec ResNet-18. For a pair of candidate concepts sets for neurons $n$ and $m$, we define their textual similarity $\mathcal{S}_{n,m}$ as the spatial mean of $L(T_{n})\cdot L(T_{m})^{\top}$ where $L(\cdot)$ is the CLIP-ViT-B/16 text encoder. A set of neurons is considered similar if $\mathcal{S}_{n,m}\geq\phi=0.8$, where $\phi$ controls the minimum similarity threshold between concepts. For practicality, we confine each neuron to exactly one group of highly similar neurons. Low-level concepts are then classified by GPT one-shot classification into the 6 NAIP superclasses.
Figure 4 presents concepts from Layer 2, color-coded by superclass. We find that clusters (each bar in Figure 4) containing more neurons are frequently associated with more interpretable features, while clusters with relatively less neurons are related to vague or irrelevant concepts. Based on these insights, we prune neurons with uninterpretable concepts in Section 5.2. We also note that "Herbaceous/Shrubland" and "Water/wetlands" are not associated with concepts in Layer 2. We hypothesize that this is likely due to the intrinsic similarity between the classes and bias in the NAIP dataset. Concepts related to "Herbaceous/Shrubland" are closely related to the "Planted/Cultivated" superclass while "Water/wetlands" images only comprises of $\sim$0.275of the dataset.
5.2 Pruning Uninterpretable Neurons
Table 6: Pruning uninterpretable neurons in Tile2Vec ResNet18.
Layer % Neurons Pruned Rand. Acc.(%) Avg. Acc. (%)
No pruning 0.00 71.63
All pruned 100.00 35.96
Layer 1 23.44 69.3 71.07
Layer 2 50.00 68.96 71.54
Layer 3 26.95 70.13 71.75
Layer 4 58.98 69.95 72.04
Layer 5 56.45 70.99 71.76
We conduct a study to identify neurons within Tile2Vec ResNet-18 model that contribute minimally to model accuracy.
Based on Section 5.1, we locate a subset of ungrouped neurons which correlate poorly to other neurons in the network. For our experiment, we define poorly correlating subsets as concepts that activate on only a single neuron. Following the original implementation, we train a classifier on embeddings from a pruned Tile2Vec ResNet-18 model.
Table 6 shows model accuracy after identifying and pruning poorly correlating neuron subsets across each layer of the model. Due to high variance in the NAIP dataset, we evaluate the baseline accuracy for a fully pruned network. In this case, we find the prediction is always the "tomatoes" subclass which achieves an accuracy of 35.96% since this subclass comprises of $\sim$35.2of the data. Our results show that a significant proportion of neurons in the model do not contribute to the overall classification. Particularly in Layers 4 and 5, we are able to prune over 50% of neurons in each layer while achieving a result better than or equal to the baseline performance (no neurons pruned). Across the entire network, the relative small difference in accuracy after pruning suggests human interpretable neurons account for more critical roles within image classification networks compared to uninterpretable neurons. Because we retrain the classifier on the pruned pipeline, the difference in accuracy using DnD pruning and that of random pruning is fairly minor, but we can see random pruning consistently suffers an accuracy drop while DnD pruning does not. In Appendix A.5, we conduct a similar experiment with a fixed classifier.
5.3 Characterizing Spurious Correlations
We characterize spurious correlations in the intermediate layers of EuroSAT-trained ResNet-50 by determining common neuron labels through Term Frequency Analysis and studying their relationship to the task. We prune these neurons to further understand the class-wise correlation of spurious concepts, shown in Table 7. Though "fishing" is the most prevalent concept in Layer 4 (41.65% of all neurons), pruning them has no impact on model accuracy. In other words, these fishing neurons are irrelevant to the classification task. "Pink" and "purple" account for 29.74% of Layer 4 neurons and have a much greater impact. These concepts, which are seemingly unrelated to the task, are spuriously correlated to the Forest, Herbaceous Vegetation, Industrial, Pasture, Residential, and Sea-Lake classes. However, the Annual Crop, Highway, and River classes have weaker correlations with these concepts.
Table 7: Pruning concepts from Layer 4 of EuroSAT-trained ResNet-50.
Concepts pruned % of neurons pruned Class-wise Accuracy (%) Avg. Acc. (%)
Ann.-Crop Forest Herb-Veget. Highway Industrial Pasture Perm.-Crop Residential River Sea-Lake
No pruning 0 95.00 98.00 93.67 96.00 96.80 91.50 90.80 98.67 92.80 97.33 95.26
Fishing 41.65 95.00 98.00 93.67 96.00 96.80 91.50 90.80 98.67 92.80 97.33 95.26
Pink/purple 29.74 88.33 0.00 0.00 62.00 0.00 0.00 18.80 0.00 76.00 0.00 24.33
@@ 6 Conclusions
In this paper, we presented Describe-and-Dissect (DnD), a novel method for automatically labeling the functionality of deep vision neurons without the need for labeled training data or a provided concept set. We accomplish this through three important steps including probing set augmentation, candidate concept generation through off-the-shelf general purpose models, and best concept selection with carefully designed scoring functions. Through extensive qualitative, quantitative, and use-case analysis, we show that DnD outperforms prior work by providing higher-quality neuron descriptions, greater generality and flexibility, and significant potential for social impact.
Acknowledgments
This work is supported in part by National Science Foundation (NSF) awards CNS-1730158, ACI-1540112, ACI-1541349, OAC-1826967, OAC-2112167, CNS-2100237, CNS-2120019, the University of California Office of the President, and the University of California San Diego’s California Institute for Telecommunications and Information Technology/Qualcomm Institute. Thanks to CENIC for the 100Gbps networks. This work used Expanse CPU, GPU and Storage at SDSC through allocation CIS230152 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support program, which is supported by National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296. The authors thank REHS program (Research Experience for High School students) in San Diego Supercomputer Center. The authors are partially supported by National Science Foundation under Grant No. 2107189, 2313105, 2430539, Hellman Fellowship, and Intel Rising Star Faculty Award. The authors would also like to thank anonymous reviewers and the editor for valuable feedback to improve the manuscript.